OpenStudy (anonymous):

Triangle ABC and Triangle A'B'C' are similar triangles with A'B' over AB = 5/4. If the area of Triangle ABC is 80 square units, find the area of Triangle A'B'C'.

OpenStudy (anonymous):

If A'B'/AB =5/4 then the Area of triangle A'B'C'/Area of ABC = 25/16 since the area must be something square therefore set x as area of A'B'C' we have x/80 =25/16 then you cross multiply to find x